OpenStudy (anonymous):

A parachutist's speed during a free fall reaches 216 kilometers per hour. What is this speed in meters per second? At this speed, how many meters will the parachutist fall during 2 seconds of free fall? Do not round your answers.

OpenStudy (anonymous):

@cwrw238 @agent0smith

OpenStudy (anonymous):

216/60=3.6

OpenStudy (anonymous):

km/h * 1000/60*60 = m/s and during free fall accel = +9.8 m/s s= ut+ 1/2 a t^2 u is initial speed s is the distance in meters

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

no idea whT um talking bout?

OpenStudy (agent0smith):

216/60=3.6 divide by 60 again to get speed in m/s

OpenStudy (anonymous):

thanks, i got this answer ^_^ i need help with others.. i'll tag you

OpenStudy (agent0smith):

k, won't be back for a while tho